# Implicit Regularization of Infinitesimally-perturbed Gradient Descent Toward Low-dimensional Solutions

### v1

@@ 6 Application: Over-parameterized Matrix Sensing
with $\alpha=\frac{1}{2}\sigma_{r}^{\star}$ and $\beta=\frac{425}{64}\sigma_{1}^{\star}$. To proceed, we first show that $\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}(\mathbf{X})\mathbf{R}_{\mathbf{X}}=\mathcal{P}_{{\mathcal{X}}^{\star}_{\leq r}}\left(\bar{\mathbf{X}}\right)$. Note that $\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}(\mathbf{X})=\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}_{\mathbf{X}}$ where $\mathbf{O}_{\mathbf{X}}$ is given by
$\mathbf{O}_{\mathbf{X}}$ $=\operatorname*{arg\,min}_{\mathbf{O}\in\mathcal{O}_{r\times r^{\prime}}}\left\lVert\mathbf{X}-\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}\right\rVert_{\mathrm{F}}$ (109)
$=\operatorname*{arg\,min}_{\mathbf{O}\in\mathcal{O}_{r\times r^{\prime}}}\left\lVert\bar{\mathbf{X}}\mathbf{R}_{\mathbf{X}}^{\top}-\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}\right\rVert_{\mathrm{F}}$
$=\operatorname*{arg\,min}_{\mathbf{O}\in\mathcal{O}_{r\times r^{\prime}}}\left\lVert\bar{\mathbf{X}}-\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}\mathbf{R}_{\mathbf{X}}\right\rVert_{\mathrm{F}}.$
@@
$\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}(\mathbf{X})\mathbf{R}_{\mathbf{X}}=\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}_{\mathbf{X}}\mathbf{R}_{\mathbf{X}}=V^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}_{\bar{\mathbf{X}}}=\mathcal{P}_{{\mathcal{X}}^{\star}_{\leq r}}\left(\bar{\mathbf{X}}\right).$
@@
$\left\langle\nabla f(\mathbf{X}),\mathbf{X}-\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}(\mathbf{X})\right\rangle$ $=\left\langle\nabla\bar{f}\left(\bar{\mathbf{X}}\right)\mathbf{R}_{\mathbf{X}}^{\top},\mathbf{X}-\mathcal{P}_{\mathcal{X}_{\mathcal{M}}}(\mathbf{X})\right\rangle$ (110)
$=\left\langle\nabla\bar{f}\left(\bar{\mathbf{X}}\right),\bar{\mathbf{X}}-\mathcal{P}_{{\mathcal{X}}^{\star}_{\leq r}}\left(\bar{\mathbf{X}}\right)\right\rangle$
@@
$=\frac{\alpha}{2}\operatorname{dist}^{2}(\mathbf{X},\mathcal{X}_{\mathcal{M}})+\frac{1}{2\beta}\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}^{2},$
where, in the last equality, we use $\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right)=\operatorname{dist}(\mathbf{X},\mathcal{X}_{\mathcal{M}})$ and $\left\lVert\nabla\bar{f}\left(\bar{\mathbf{X}}\right)\right\rVert_{\mathrm{F}}=\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}$, established in Equation 105 and Equation 106, respectively. This completes the proof. ∎
@@
$R(\tau)=\sup_{\mathbf{X}\in\mathcal{N}_{\mathcal{M}}(\tau)}r(\mathbf{X})\leq\sup_{\mathbf{X}\in\mathcal{N}_{\mathcal{M}}(\tau)}3r_{+}(\mathbf{X})=\sup_{\mathbf{X}\in\mathcal{N}_{\mathcal{M}}(\tau)-\mathcal{M}}\left\{-\frac{3\nabla^{2}_{-}f\big{(}\mathbf{X}^{\sharp}\big{)}\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]}{\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}}\right\},$ (111)
where the first inequality is due to the fact that $r_{-}(\mathbf{X})\leq 0$ by definition. Therefore, to control $R(\tau)$, it suffices to provide a uniform upper bound on $-\nabla^{2}_{-}f\big{(}\mathbf{X}^{\sharp}\big{)}\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]$ for every $\mathbf{X}\in\mathcal{N}_{\mathcal{M}}(\tau)-\mathcal{M}$. Recall that
@@
Here $\nabla^{2}F(\mathbf{X}^{\sharp})$ is the expected version of $\nabla^{2}f(\mathbf{X}^{\sharp})$ (after setting $\delta=0$ in $(\delta,r)$-RIP) and $\nabla^{2}H(\mathbf{X}^{\sharp})$ is the positive semi-definite part of $\nabla^{2}F(\mathbf{X}^{\sharp})$. Now, we are ready to control $-\nabla^{2}_{-}f\big{(}\mathbf{X}^{\sharp}\big{)}\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]$. First, we have
@@
$\leq\sqrt{r}\delta\max\left\{\sigma_{1}^{\star},\left\lVert\mathbf{X}^{\sharp}\right\rVert^{2}\right\}+2r\delta\left\lVert\mathbf{X}^{\sharp}\right\rVert^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}12r\sigma_{1}^{\star}\delta,$
where, in $(a)$, we use $(\delta,2r)$-RIP, and in $(b)$ we use the fact that $\left\lVert\mathbf{X}^{\sharp}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}$. Next, we control $-\nabla^{2}_{-}F\big{(}\mathbf{X}^{\sharp}\big{)}\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]$. Evidently, we have $\nabla^{2}H(\mathbf{X}^{\sharp})\succeq\bm{0}$. This implies that
$-\nabla^{2}_{-}F\big{(}\mathbf{X}^{\sharp}\big{)}\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]\leq-\nabla^{2}_{-}(F-H)\big{(}\mathbf{X}^{\sharp}\big{)}\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]=\left\langle{\mathbf{M}^{\star}}\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right\rangle,$ (117)
where the equality holds since $\nabla^{2}(F-H)\big{(}\mathbf{X}^{\sharp}\big{)}\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]=-\left\langle{\mathbf{M}^{\star}}\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right\rangle\leq 0$, which in turn implies that $\nabla^{2}(F-H)\big{(}\mathbf{X}^{\sharp}\big{)}=\nabla^{2}_{-}(F-H)\big{(}\mathbf{X}^{\sharp}\big{)}$. Furthermore, we have
$\left\langle{\mathbf{M}^{\star}}\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right\rangle$ $\leq\left\lVert{\mathbf{M}^{\star}}\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}\leq\sigma_{1}^{\star}\left\lVert\mathcal{P}_{\mathbf{V}^{\star}}\mathcal{P}_{\mathbf{X}^{\perp}}\right\rVert\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}\leq 3\sigma_{1}^{\star}\sqrt{r}\kappa\delta\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}.$ (118)
In the last inequality, we use the fact that $\left\lVert\mathcal{P}_{\mathbf{V}^{\star}}\mathcal{P}_{\mathbf{X}^{\perp}}\right\rVert=\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}^{\sharp}}\right\rVert\leq 4\sqrt{r}\kappa\delta$ since $\mathbf{X}^{\sharp}\in\mathcal{M}$. Combining the obtained bounds in Equation 118 and Equation 116 with Equation 115, we obtain
$-\nabla^{2}_{-}f\big{(}\mathbf{X}^{\sharp}\big{)}\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]\leq\left(3\sigma_{1}^{\star}\sqrt{r}\kappa\delta+12\sigma_{1}^{\star}r\delta\right)\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}\leq 15\sigma_{1}^{\star}r\kappa\delta\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}.$ (119)
Combining this inequality with Equation 111 yields $R(\tau)\leq 45\sigma_{1}^{\star}\kappa r\delta$ with $\tau=\frac{1}{10}\sqrt{\frac{\sigma_{1}^{\star}}{d}}$. ∎
@@
With all the conditions in place, we are now ready to establish the near-linear convergence of IPGD+ for the over-parameterized matrix sensing.
@@
The proof is an application of Theorem 8. Indeed, Propositions 23, 24, 28 and 30 hold under the conditions of Theorem 31. Specifically, we have
  • $f(\mathbf{X})$ in (MS) is ${(L,\mathcal{M},\tau)}$-gradient-Lipschitz and ${(\rho,\mathcal{M},\tau)}$-Hessian-Lipschitz with parameters $L=15\sigma_{1}^{\star}$, $\rho=15\sqrt{\sigma_{1}^{\star}}$, and $\tau\leq\frac{1}{10}\sqrt{\frac{\sigma_{1}^{\star}}{d}}$.
@@
  • $\mathcal{M}$ satisfies $\left(\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}}\right)$$-\mathcal{M}$-strict saddle property with parameters $\bar{\epsilon}_{\mathrm{g}}=\frac{1}{160}\sigma_{r}^{\star 3/2},\bar{\epsilon}_{\mathrm{H}}=\frac{1}{5}\sigma_{r}^{\star},\bar{\epsilon}_{\mathcal{M}}=\frac{1}{8}\sigma_{r}^{\star 1/2}$. Moreover, it satisfies $\left(\alpha,\beta,\zeta\right)$$-\mathcal{M}$-regularity property with parameters $\alpha=\frac{1}{2}\sigma_{r}^{\star},\beta=\frac{425}{64}\sigma_{1}^{\star},\zeta=\frac{1}{4}\sigma_{r}^{\star 1/2}$.
@@
  • Due to our choice of the initial point and the definition of the implicit region in Equation 78, we have $\left\lVert\mathbf{X}_{0}^{\perp}\right\rVert=0$. Moreover, due to RIP, we have $\Delta_{f}=f(\mathbf{X}_{0})-f^{\star}=\frac{1}{4}\left\lVert\mathcal{A}({\mathbf{M}^{\star}})\right\rVert^{2}\leq\frac{1+\delta}{4}\left\lVert{\mathbf{M}^{\star}}\right\rVert_{\mathrm{F}}^{2}\leq\frac{r}{2}\sigma_{1}^{\star 2}$.
@@
iterations, IPGD+ outputs a solution $\mathbf{X}_{T}$ satisfying $\operatorname{dist}(\mathbf{X}_{T},\mathcal{X}_{\mathcal{M}})\leq\frac{\epsilon}{15\sigma_{1}^{\star}}$. Here, we set $\delta=O\Big{(}\frac{1}{r^{2}\kappa^{5}\left(\log^{4}(1/\gamma)+\log^{4}(\sigma_{1}^{\star}d/\chi)\right)}\Big{)}$ to ensure that $R(\tau)=O(1/(\eta T))$, as required by Theorem 8. Moreover, since $\mathcal{X}_{\mathcal{M}}=\mathcal{X}^{\star}$ when $\delta\leq\frac{1}{10}$, it follows that $\operatorname{dist}(\mathbf{X}_{T},\mathcal{X}^{\star})\leq\frac{\epsilon}{15\sigma_{1}^{\star}}$, which completes the proof. ∎
Next, we present a few key observations based on Theorem 31.
@@
To apply IPGD+ to the over-parameterized matrix sensing problem, we initialize the algorithm at the origin. It is straightforward to verify that the origin is a strict saddle point of (MS). As a result, IPGD+ automatically applies a small random perturbation to the initial point, allowing the algorithm to escape this saddle point. This strategy is akin to the small random initialization used for GD in [6].
@@
The convergence rate of IPGD+ toward the true rank-$r$ solution scales as $O(\log^{4}(1/\epsilon))$. While this convergence rate is significantly better than that of general saddle—escaping methods such as PGD—which fail to converge to the true solution in the over-parameterized setting where $r^{\prime}>r$—it remains worse than the linear convergence rate of $O(\log(1/\epsilon))$ achieved by GD with small random initialization [6]. This suboptimality primarily arises because our algorithm is designed to handle the worst-case landscape, where it may encounter up to $O(\log^{3}(1/\epsilon))$ saddle points. In contrast, [60] shows that for over-parameterized matrix sensing, GD encounters at most $r$ saddle points before reaching the true rank-$r$ solution. We believe this insight can also be incorporated into the analysis of IPGD+ to further improve its convergence rate.

### v2

@@ 6 Application: Over-parameterized Matrix Sensing
with $\alpha=\frac{1}{2}\sigma_{r}^{\star}$ and $\beta=\frac{425}{64}\sigma_{1}^{\star}$. To proceed, we first show that $\mathcal{P}_{\mathcal{X}_{\mathcal{M}_{\text{RIP}}}}(\mathbf{X})\mathbf{R}_{\mathbf{X}}=\mathcal{P}_{{\mathcal{X}}^{\star}_{\leq r}}\left(\bar{\mathbf{X}}\right)$. Note that $\mathcal{P}_{\mathcal{X}_{\mathcal{M}_{\text{RIP}}}}(\mathbf{X})=\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}_{\mathbf{X}}$ where $\mathbf{O}_{\mathbf{X}}$ is given by
$\mathbf{O}_{\mathbf{X}}$ $=\operatorname*{arg\,min}_{\mathbf{O}\in\mathcal{O}_{r\times r^{\prime}}}\left\lVert\mathbf{X}-\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}\right\rVert_{\mathrm{F}}=\operatorname*{arg\,min}_{\mathbf{O}\in\mathcal{O}_{r\times r^{\prime}}}\left\lVert\bar{\mathbf{X}}\mathbf{R}_{\mathbf{X}}^{\top}-\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}\right\rVert_{\mathrm{F}}=\operatorname*{arg\,min}_{\mathbf{O}\in\mathcal{O}_{r\times r^{\prime}}}\left\lVert\bar{\mathbf{X}}-\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}\mathbf{R}_{\mathbf{X}}\right\rVert_{\mathrm{F}}.$ (110)
@@
$\mathcal{P}_{\mathcal{X}_{\mathcal{M}_{\text{RIP}}}}(\mathbf{X})\mathbf{R}_{\mathbf{X}}=\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}_{\mathbf{X}}\mathbf{R}_{\mathbf{X}}=V^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}_{\bar{\mathbf{X}}}=\mathcal{P}_{{\mathcal{X}}^{\star}_{\leq r}}\left(\bar{\mathbf{X}}\right).$
@@
$\left\langle\nabla f(\mathbf{X}),\mathbf{X}-\mathcal{P}_{\mathcal{X}_{\mathcal{M}_{\text{RIP}}}}(\mathbf{X})\right\rangle$ $=\left\langle\nabla\bar{f}\left(\bar{\mathbf{X}}\right)\mathbf{R}_{\mathbf{X}}^{\top},\mathbf{X}-\mathcal{P}_{\mathcal{X}_{\mathcal{M}_{\text{RIP}}}}(\mathbf{X})\right\rangle=\left\langle\nabla\bar{f}\left(\bar{\mathbf{X}}\right),\bar{\mathbf{X}}-\mathcal{P}_{{\mathcal{X}}^{\star}_{\leq r}}\left(\bar{\mathbf{X}}\right)\right\rangle$ (111)
@@
$=\frac{\alpha}{2}\operatorname{dist}^{2}(\mathbf{X},\mathcal{X}_{\mathcal{M}_{\text{RIP}}})+\frac{1}{2\beta}\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}^{2},$
where, in the last equality, we use $\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right)=\operatorname{dist}(\mathbf{X},\mathcal{X}_{\mathcal{M}_{\text{RIP}}})$ and $\left\lVert\nabla\bar{f}\left(\bar{\mathbf{X}}\right)\right\rVert_{\mathrm{F}}=\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}$, established in Equation 106 and Equation 107, respectively. This completes the proof. ∎
@@
$R(\tau)=\sup_{\mathbf{X}\in\mathcal{N}_{\mathcal{M}_{\text{RIP}}}(\tau)}r(\mathbf{X})\leq\sup_{\mathbf{X}\in\mathcal{N}_{\mathcal{M}_{\text{RIP}}}(\tau)}3r_{+}(\mathbf{X})=\sup_{\mathbf{X}\in\mathcal{N}_{\mathcal{M}_{\text{RIP}}}(\tau)-\mathcal{M}_{\text{RIP}}}\left\{-\frac{3\nabla^{2}_{-}f\big(\mathbf{X}^{\sharp}\big)\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]}{\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}}\right\},$ (112)
where the first inequality is due to the fact that $r_{-}(\mathbf{X})\leq 0$ by definition. Therefore, to control $R(\tau)$, it suffices to provide a uniform upper bound on $\nabla^{2}_{-}f\big(\mathbf{X}^{\sharp}\big)\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]$ for every $\mathbf{X}\in\mathcal{N}_{\mathcal{M}_{\text{RIP}}}(\tau)-\mathcal{M}_{\text{RIP}}$. Recall that
@@
Here $\nabla^{2}F(\mathbf{X}^{\sharp})$ is the expected version of $\nabla^{2}f(\mathbf{X}^{\sharp})$ (after setting $\delta=0$ in $(\delta,r)$-RIP) and $\nabla^{2}H(\mathbf{X}^{\sharp})$ is the positive semi-definite part of $\nabla^{2}F(\mathbf{X}^{\sharp})$. Now, we are ready to control $\nabla^{2}_{-}f\big(\mathbf{X}^{\sharp}\big)\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]$. First, we have
@@
[…]
where, in $(a)$, we use $(\delta,2r)$-RIP, and in $(b)$ we use the fact that $\left\lVert\mathbf{X}^{\sharp}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}$. Next, we control $\nabla^{2}_{-}F\big(\mathbf{X}^{\sharp}\big)\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]$. Evidently, we have $\nabla^{2}H(\mathbf{X}^{\sharp})\succeq\bm{0}$. This implies that
$-\nabla^{2}_{-}F\big(\mathbf{X}^{\sharp}\big)\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]\leq-\nabla^{2}_{-}(F-H)\big(\mathbf{X}^{\sharp}\big)\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]=\left\langle\mathbf{\Theta}^{\star}\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right\rangle,$ (118)
where the equality holds since $\nabla^{2}(F-H)\big(\mathbf{X}^{\sharp}\big)\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]=-\left\langle\mathbf{\Theta}^{\star}\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right\rangle\leq 0$, which in turn implies that $\nabla^{2}(F-H)\big(\mathbf{X}^{\sharp}\big)=\nabla^{2}_{-}(F-H)\big(\mathbf{X}^{\sharp}\big)$. Furthermore, we have
$\left\langle\mathbf{\Theta}^{\star}\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right\rangle$ $\leq\left\lVert\mathbf{\Theta}^{\star}\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}\leq\sigma_{1}^{\star}\left\lVert\mathcal{P}_{\mathbf{V}^{\star}}\mathcal{P}_{\mathbf{X}^{\perp}}\right\rVert\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}\leq 3\sigma_{1}^{\star}\sqrt{r}\kappa\delta\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}.$ (119)
In the last inequality, we use $\left\lVert\mathcal{P}_{\mathbf{V}^{\star}}\mathcal{P}_{\mathbf{X}^{\perp}}\right\rVert=\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}^{\sharp}}\right\rVert\leq 4\sqrt{r}\kappa\delta$ since $\mathbf{X}^{\sharp}\in\mathcal{M}_{\text{RIP}}$. Combining the obtained bounds in Equation 119 and Equation 117 with Equation 116, we obtain
$\nabla^{2}_{-}f\big(\mathbf{X}^{\sharp}\big)\left[\mathbf{X}^{\perp},\mathbf{X}^{\perp}\right]\leq\left(3\sigma_{1}^{\star}\sqrt{r}\kappa\delta+12\sigma_{1}^{\star}r\delta\right)\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}\leq 15\sigma_{1}^{\star}r\kappa\delta\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}^{2}.$ (120)
Combining this inequality with Equation 112 yields $R(\tau)\leq 45\sigma_{1}^{\star}\kappa r\delta$ with $\tau=\frac{1}{10}\sqrt{\frac{\sigma_{1}^{\star}}{d}}$.
@@
With all the conditions in place, we are ready to prove the convergence of IPGD+ for the over-parameterized matrix sensing.
@@
The proof is an application of Theorem 4. Indeed, Propositions 6, 7, 8 and 9 hold under the conditions of Theorem 5. Specifically, we have
  • $f(\mathbf{X})$ in (MS) is $(L,\mathcal{M}_{\text{RIP}},\tau)$}-gradient-Lipschitz and $(\rho,\mathcal{M}_{\text{RIP}},\tau)$}-Hessian-Lipschitz with parameters $L=15\sigma_{1}^{\star}$, $\rho=15\sqrt{\sigma_{1}^{\star}}$, and $\tau\leq\frac{1}{10}\sqrt{\frac{\sigma_{1}^{\star}}{d}}$.
@@
  • $\mathcal{M}_{\text{RIP}}$ satisfies $\left(\bar{\epsilon}_{\mathrm{g}},\bar{\epsilon}_{\mathrm{H}},\bar{\epsilon}_{\mathcal{M}}\right)$-$\mathcal{M}_{\text{RIP}}$-strict saddle property with parameters $\bar{\epsilon}_{\mathrm{g}}=\frac{1}{160}\sigma_{r}^{\star 3/2},\bar{\epsilon}_{\mathrm{H}}=\frac{1}{5}\sigma_{r}^{\star},\bar{\epsilon}_{\mathcal{M}}=\frac{1}{8}\sigma_{r}^{\star 1/2}$. Moreover, it satisfies $\left(\alpha,\beta,\zeta\right)$-$\mathcal{M}_{\text{RIP}}$-regularity property with parameters $\alpha=\frac{1}{2}\sigma_{r}^{\star},\beta=\frac{425}{64}\sigma_{1}^{\star},\zeta=\frac{1}{4}\sigma_{r}^{\star 1/2}$.
@@
  • Due to our choice of the initial point and the definition of the implicit region in Equation 79, we have $\left\lVert\mathbf{X}_{0}^{\perp}\right\rVert=0$. Moreover, due to RIP, we have $\Delta_{f}=f(\mathbf{X}_{0})-f^{\star}=\frac{1}{4}\left\lVert\mathcal{A}(\mathbf{\Theta}^{\star})\right\rVert^{2}\leq\frac{1+\delta}{4}\left\lVert\mathbf{\Theta}^{\star}\right\rVert_{\mathrm{F}}^{2}\leq\frac{r}{2}\sigma_{1}^{\star 2}$.
@@
iterations, IPGD+ outputs a solution $\mathbf{X}_{T}$ satisfying $\operatorname{dist}(\mathbf{X}_{T},\mathcal{X}_{\mathcal{M}_{\text{RIP}}})\leq\frac{\epsilon}{15\sigma_{1}^{\star}}$. Here, we set $\delta=O\Big(\frac{1}{r^{2}\kappa^{5}\left(\log^{4}(1/\gamma)+\log^{4}(\sigma_{1}^{\star}d/\chi)\right)}\Big)$ to ensure that $R(\tau)=O(1/(\eta T))$, as required by Theorem 4. Moreover, since $\mathcal{X}_{\mathcal{M}_{\text{RIP}}}=\mathcal{X}^{\star}$ when $\delta\leq\frac{1}{10}$, it follows that $\operatorname{dist}(\mathbf{X}_{T},\mathcal{X}^{\star})\leq\frac{\epsilon}{15\sigma_{1}^{\star}}$.
@@
To apply IPGD+ to the over-parameterized matrix sensing problem, we initialize the algorithm at the origin. It is straightforward to verify that the origin belongs to $\mathcal{M}_{\text{RIP}}$ and is a strict saddle point of (MS). As a result, IPGD+ automatically applies a small random perturbation to the initial point, allowing the algorithm to escape this saddle point. This strategy is akin to the small random initialization used for GD in [stoger2021small].
@@
The convergence rate of IPGD+ toward the true rank-$r$ solution scales as $O(\log^{4}(1/\epsilon))$. While this convergence rate is significantly better than that of general saddle—escaping methods such as PGD—which fail to converge to the true solution in the over-parameterized setting where $r^{\prime}>r$—it remains worse than the linear convergence rate of $O(\log(1/\epsilon))$ achieved by GD with small random initialization [stoger2021small]. This suboptimality primarily arises because our algorithm is designed to handle the worst-case landscape, where it may encounter up to $O(\log^{3}(1/\epsilon))$ saddle points. In contrast, [jin2023understanding] shows that for over-parameterized matrix sensing, GD encounters at most $r$ saddle points before reaching the true rank-$r$ solution. We believe this insight can also be incorporated into the analysis of IPGD+ to further improve its convergence rate.